OpenStudy (anonymous):

Solve the equation for y: 2x + 3(y + 9) = 17

jimthompson5910 (jim_thompson5910):

I'll get you started 2x + 3(y + 9) = 17 2x + 3y + 27 = 17 2x + 3y = 17-27 2x + 3y = -10